OpenStudy (anonymous):

i need help wth this question If 6x - 3y = 30 and 4x = 2 - y then find x + y.

OpenStudy (anonymous):

can someone please help me

OpenStudy (anonymous):

Get the second equation in y=mx+b format (4x=2-y) add y to both sides (4x+y=2) subtract 4x from both sides (y=-4x-2) Plug this into the first equation {6x-3*(-4x+2)=30} distribute the -3 (6x+12x-6=30) Add 6 to both sides (6x+12x=30+6) Combine like terms (18x=36) divide by 18 x=2 Plug this back into the second equation (4*(2)=2-y) (8=2-y) subtract 2 from both sides (6=-y) negative variables is a no no .... change all the signs y=-6 Plug x and y into x+y (2+(-6)) -4